OpenStudy (anonymous):

Smart people please help! True or False question! The polar coordinates (1, 0) expressed as Cartesian coordinates are (0, 1), T/F? Please help me!

OpenStudy (anonymous):

\[ \rho =1\\ \theta =0\\ x= \rho \cos(\theta)= 1\\ y =\rho \sin(\theta)= 0 \] False it should be (1,0) in cartesian coordinates

OpenStudy (anonymous):